Tracy needs to find the length and width of the equipment storage room where she works. The room is 1 meter longer then it is wide. The perimeter is 14 meters
Answer by graphmatics(170)

Let x be how wide the room is. Then x+1 is how long the room is. The perimeter of a rectangle (the room) is 2*x + 2*(x+1) so
2*x + 2*(x+1) = 14
2*x + 2*x + 2 = 14
2*x + 2*x = 14 - 2
4*x = 12
x = 3
So the room is 3 meters wide and 4 meters long.